write the equation of the line wuth x intercept (12,0) and y intercept (0,-2). Thank you.
Answer by checkley71(8403)

slope=(-2-0)/(0-12)=-2/-12=1/6.
WITH A Y INTERCEPT OF -2 WE HAVE THE LINE FORMULA
Y=X/6-2